OpenStudy (anonymous):

solve using the elimination method -7x - y=-30

OpenStudy (asnaseer):

you have two unknowns here: x and y to use the elimination method you therefore need two equations. you have only specified one equation.

OpenStudy (anonymous):

4x + y =21

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

We need one more equ. for using elimination meth.

OpenStudy (asnaseer):

so the two equations you have are: -7x - y = -30 4x + y = 21

OpenStudy (asnaseer):

can you see that if you add these two equations, then the terms involving y will be eliminated?

OpenStudy (anonymous):

Yes I think so...

OpenStudy (asnaseer):

so what do you get if you add these two equations?

OpenStudy (anonymous):

we do that for eliminating +y and -y

OpenStudy (asnaseer):

@warsop87 please ask questions if you do not understand any specific point.

OpenStudy (anonymous):

I got -30

OpenStudy (anonymous):

I think im doing this all wrong

OpenStudy (asnaseer):